A positron is an elementary particle identical to an electron except that its charge is +e. An electron and a positron can rotate about their center of mass as if they were a dumbbell connected by a massless rod. Part A What is the orbital frequency in rev/s for an electron and a positron 1.30 nm apart? I got 5.1*10^13 and it is wrong.
Answer #1

Force, F = mv^2/(d/2) = kq^2/d^2

V^2 = kq^2/2 md

V^2 = (9 x 10^9 x (1.6 x 10^-19)^2)/(2 x 9.1 x 10^-31 x 1.3 x 10^-9)

V = 3.12 x 10^5 m/s

Distance, d = 2 pi x (1.3 x 10^-9)/2 = 4.08 x 10^-9 m

Time, t = d/v = 1.31 x 10^-14 sec

Frequency = 1/t = 7.64 x 10^13 Hz

Comment in case any doubt please rate my answer ....